Eye Trendy Corporation is a distributor of frames for sunglasses. The company’s controller is currently preparing a budget for the third quarter of the year. The following information is from company’s financial records: Projected Sales July 3,120 units August 2,000 units September 2,640 units October 3,000 units • Selling price is RM25 per unit • Collections from customers are normally 70 per cent in the month of sale, 20 per cent in the month following sale, and 9 per cent in the second month following the sale. The balance is expected to be uncollectible. Projected Purchases • Purchase price is RM18 per unit. • All frames purchases are on account. 70 per cent of the frames purchased are paid for in the month of purchase; the remaining 30 per cent are paid for in the month after acquisition. • Inventory of frames on 1st July is 1,200 units. The frames inventory at the end of each month equals 20 per cent of sales anticipated for the following month. • The company purchases the frames as needed in multiple quantities of 1,000 units per shipment. Operating Expenses • General and administrative expenses are projected to be RM33,000 for the quarter. The breakdown of these expenses is presented in the following schedule. All cash expenditures will be paid uniformly throughout the quarter: Promotion RM9,000 Insurance RM12,000 Utilities RM7,500 Depreciation RM4,500 Total RM33,000 Other information • Cash proceeds from sale of old equipment amounted to RM5,000 in the month of August. • Purchase of new equipment amounted to RM50,000 is to be made in the month of September. • Eye Trendy is expected to maintain a minimum cash balance of RM20,000 at all times. If the cash balance is less than RM20,000 at the end of each month, the company borrows amounts necessary to maintain this balance. All amounts are repaid out of the subsequent positive cash flow. • The company’s cash balance on 1st July is RM22,000. Required: a. Prepare the following schedules: (i) Expected cash collections for the sales of frames during the third quarter. Show computations by month and in total for the quarter. (7 marks) (ii) Expected Cash disbursements for the purchases of frames during the third quarter. Show computations by month and in total for the quarter. (8 marks) (iii) Expected Cash balance on 30th September. Show computations by month and in total for the quarter. Answer with different person because previous answer has each different answer.
Answer #1
Please give positive ratings so I can keep answering. If you have any queries please comment. Thanks!
Eye Trendy Corporation Amount RM
Sales Budget July August September October Total Jul- Sep Qtr. Formula
Sales units            3,120.00      2,000.00        2,640.00      3,000.00                7,760.00 A
Selling Price                 25.00           25.00             25.00           25.00 B
Budgeted Sales Revenue         78,000.00 50,000.00     66,000.00 75,000.00            194,000.00 C=A*B
Collection Budget July August September October Total Jul- Sep Qtr.
70 % in the month of sale         54,600.00 35,000.00      46,200.00 52,500.00            135,800.00 D=C*70%
20 % in the month following sale 15,600.00      10,000.00 13,200.00              25,600.00 E= 20% of C of previous month
9 % in the second month following the sale        7,020.00      4,500.00                7,020.00 F= 9% of C of previous to previous month
Total Scheduled Collections         54,600.00 50,600.00     63,220.00 70,200.00            168,420.00 G=D+E+F
Inventory Purchase Budget July August September October Total Jul- Sep Qtr.
Sales units            3,120.00      2,000.00        2,640.00      3,000.00                7,760.00 A
Closing Stock @ 20%               400.00         528.00           600.00 H= 20% of A of next month
Opening Stock            1,200.00         400.00           528.00 I=H of previous month, July figure is given in sum.
Purchase required           2,320.00     2,128.00        2,712.00                7,160.00 J=A+H-I
But it should be done in multiple of 1000
Budget Purchase units            2,400.00      2,200.00        2,800.00                7,400.00 K
Cost per unit                 18.00           18.00             18.00 L
Material to be purchased (Value)         43,200.00 39,600.00     50,400.00            133,200.00 M=K*L
Inventory Payment Budget July August September Total Jul- Sep Qtr.
Material to be purchased (Value)         43,200.00 39,600.00      50,400.00            133,200.00 M
Paid in current month 70%         30,240.00 27,720.00      35,280.00              93,240.00 N=M*70%
Paid in next month 30% 12,960.00      11,880.00              24,840.00 O= 30% of M of previous month.
Total Scheduled payments         30,240.00 40,680.00     47,160.00            118,080.00 P=M+O
Cash budget July August September Total Jul- Sep Qtr.
Beginning Cash Balance         22,000.00 20,000.00      20,000.00
Plus: Collections from sales         54,600.00 50,600.00      63,220.00            168,420.00 G
Sale of equipment      5,000.00                5,000.00
Cash Available         76,600.00 75,600.00     83,220.00
Disbursements
Inventory Payment Budget         30,240.00 40,680.00      47,160.00            118,080.00 P
Promotion            9,000.00      9,000.00        9,000.00              27,000.00
Insurance         12,000.00 12,000.00      12,000.00              36,000.00
Utilities            7,500.00      7,500.00        7,500.00              22,500.00
Purchase of Equipment                  -        50,000.00              50,000.00
Total cash payments         58,740.00 69,180.00 125,660.00            253,580.00
Excess cash available over cash payments         17,860.00     6,420.00 (42,440.00)
Receipt Short Term Loan            2,140.00 13,580.00      62,440.00              78,160.00
Repayment Short Term Loan                        -                    -                      -                               -  
Ending cash balance         20,000.00 20,000.00     20,000.00
Loan Balance July August September Total Jul- Sep Qtr.
Opening Balance                        -        2,140.00      15,720.00
Receipt Short Term Loan            2,140.00 13,580.00      62,440.00              78,160.00
Repayment Short Term Loan                  -                      -  
Closing Balance           2,140.00 15,720.00     78,160.00